OpenStudy (anonymous):

Change the Cartesian Integral into an equivalent polar integral then evaluate: (int(int(dydx)) from -1 to 1 and 0 to sqrt(1-x^2)

OpenStudy (anonymous):

nvm got it...

OpenStudy (anonymous):

How'd you do it? If I may

OpenStudy (anonymous):

YOu have to graph the boundaries then find the new boundaries: 0 to pi/2 and 0 to 1. Then you use the equation int(int(rdrdtheta) which = pi/2

OpenStudy (anonymous):

limits for r are 0 to 1, theta 0 to pi

OpenStudy (anonymous):

I got 0 to pi /2 because its y...

OpenStudy (anonymous):

\[\int\limits_{}{}\int\limits_{}{}dxdy=\int\limits_{}{}\int\limits_{}{}|\frac{\partial J(x,y)}{\partial (r, \theta)}|dr d \theta=\int\limits_{}{}\int\limits_{}{}r dr d \theta\]

OpenStudy (anonymous):

Integrate out r from 0 to 1, then theta from 0 to pi.

OpenStudy (anonymous):

You get \[\frac{\pi}{2}\]

OpenStudy (anonymous):

ohh i see that thanks

OpenStudy (anonymous):

welcome. i stopped on your question and went for a shower - thought someone was taking it.